OpenStudy (anonymous):

Please help to solve 2X^3-5X^2-46X+24=0

OpenStudy (anonymous):

<=> 2X^3 - 12X^2 + 7X^2 - 42X - 4X + 24 = 0 <=> (2X^2 + 7X - 4)(X-6)=0 <=>(2X^2 + 8X - X - 4)(X-6)=0 <=>(X+4)(2X-1)(X-6)=0 Is it right?

OpenStudy (anonymous):

how come to do the 1st line and the 3rd line?

OpenStudy (anonymous):

nvrmind.....thank you. i figured out what u did :)